# Unsupervised Multimodal Graph-based Model for Geo-social Analysis

### v1

@@ 3.1 Network Architectures
A schematic overview of both architectures is provided in Fig. 1. We employ Graph Convolutional Networks (GCNs) to model local context and neighborhood interactions across semantic and geographic dimensions in both architectures. GCNs operate via a localized message-passing scheme, updating node features according to:
@@
where $\hat{A}$ is the normalized adjacency matrix, $\mathbf{X}^{(l)}$ denotes the feature matrix at layer $l$, $\mathbf{W}^{(l)}$ is a learnable weight matrix, and $\sigma$ is a non-linear activation function such as ReLU. Our implementation uses two GCN layers with ReLU activations. In the MultiGraph setup, dropout and layer normalization are applied between layers to enhance generalization and training stability. The MonoGraph architecture processes the unified graph with a shared GCN encoder to learn joint node embeddings. Conversely, the MultiGraph architecture builds two separate GCN branches, one for semantic and one for geographic graphs, producing semantic embeddings $\mathbf{x}_{i}^{(s)}$ and geographic embeddings $\mathbf{x}_{i}^{(g)}$ for each node $i$. These embeddings are then fused through a learnable multi-head attention mechanism. In this mechanism, the geographic embeddings serve as the queries $(Q)$, while the semantic embeddings act as both the keys $(K)$ and values $(V)$, enabling the model to selectively attend to relevant semantic features conditioned on geographic context:
$\mathbf{Z}_{i}=\text{Multi-Head}\big(\mathbf{Q}=\mathbf{x}_{i}^{(g)},\mathbf{K}=\mathbf{x}_{i}^{(s)},\mathbf{V}=\mathbf{x}_{i}^{(s)}\big),\quad\mathbf{z}_{i}=\frac{1}{h}\sum_{j=1}^{h}\mathbf{Z}_{i}^{(j)},$
where the multi-head attention layer first linearly projects $(Q,K,V)$ into $h$ subspaces (attention heads) of dimension $d_{k}=\frac{\text{embed\_dim}}{h}$, then computes scaled dot-product attention for each head as
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$
and finally concatenates the outputs from all attention heads into a single vector, which is then linearly transformed to produce the fused embedding. Here, $h$ denotes the number of attention heads. This fusion mechanism enables adaptive balancing of semantic and spatial information, selectively modulating the influence of nodes and edges to optimize the overall training objective.
@@ 3.2 Loss Design and Training Objective
To guide the learning of node representations, we use a composite loss that combines contrastive, coherence, and alignment objectives. The contrastive loss helps preserve local semantic structure. The coherence loss promotes cluster structure. The alignment loss further stabilizes training by encouraging each node to stay close to the centroid of its assigned cluster. By jointly optimizing these losses, the model is able to learn embeddings that capture both local neighborhood structure and global cluster organization, while also maintaining alignment with cluster prototypes. Clustering is performed using the K-means algorithm with a fixed random seed and fixed cluster numbers to ensure consistency across training epochs, although any other clustering algorithm, differentiable or not, can be used without modifying the core framework.
@@
Coherence Loss: The coherence loss consists of two components: intra-cluster similarity, which ensures topics are internally consistent, and inter-cluster dissimilarity, which discourages overlap between topics.
Intra-cluster Similarity: The intra-cluster similarity measures how similar the embeddings of nodes within the same cluster are to each other. Intuitively, nodes that belong to the same community should have similar representations, as they share common structural or semantic properties. This similarity is commonly quantified using normalized pairwise cosine similarity, which has been shown to be an effective measure for assessing the compactness of clusters, particularly for text and high-dimensional data [ghosh2006]. The cosine similarity captures the angular distance between embeddings, providing a metric that is robust to differences in magnitude while emphasizing the relative similarity of node features:
$\mu_{\text{intra}}=\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)=C(j)\}}}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}$
@@
where $K$ is the number of clusters, and $C_{k}$ denotes the set of nodes in cluster $k$. The denominator normalizes the similarity by the number of node pairs in each cluster. Cosine similarity is computed as:
$\cos(\mathbf{z}_{i},\mathbf{z}_{j})=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|_{2}\|\mathbf{z}_{j}\|_{2}}$
Here, $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ represent the feature embeddings of nodes $i$ and $j$, respectively, and are normalized by their L2 norms to ensure that only the direction (rather than magnitude) of the vectors influences the similarity measure.
Maximizing intra-cluster similarity encourages the model to learn node embeddings that are close to one another within the same cluster. This promotes compactness and internal consistency in the embedding space, improving the semantic or spatial cohesion of discovered clusters. In this context, higher values of intra-cluster similarity, typically close to 1, indicate tight and cohesive clusters, while values near 0 or -1 suggest poorly defined clusters.
Inter-cluster Similarity: The inter-cluster similarity captures how similar node embeddings are across different clusters. Ideally, embeddings from separate clusters should be dissimilar. We define this as:
$\mu_{\text{inter}}=\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)\neq C(j)\}}}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}$
Here, $P_{\text{inter}}$ denotes all node pairs from distinct clusters. High inter-cluster similarity values would indicate insufficient separation between clusters, while values close to -1 reflect distinct and well-separated clusters. So, ideally we aim for low inter-cluster similarity, with values closer to -1 indicating optimal separation between clusters.
To jointly optimize intra-cluster compactness and inter-cluster separation, we define the coherence loss:
@@
$\mathcal{L}_{\text{coherence}}=-\left(\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}-\lambda\cdot\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}\right)$
where $\lambda$ is a hyperparameter that controls the trade-off between maximizing intra-cluster similarity and minimizing inter-cluster similarity. This loss function combines both clustering principles into a single differentiable objective. The negative sign transforms the coherence score into a minimization problem, consistent with standard optimization procedures. Minimizing $\mathcal{L}_{\text{coherence}}$ leads to embedding spaces where nodes within the same cluster are tightly grouped, while nodes from different clusters are well-separated, enhancing both interpretability and performance of downstream clustering tasks.
@@
Contrastive Loss: To preserve local semantic or structural neighborhoods, we use a contrastive loss that pulls embeddings of similar nodes closer and pushes dissimilar ones apart. For a batch of positive and negative node pairs, the contrastive objective is defined as:
$\mathcal{L}_{\text{contrast}}=-\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{(i,k)\in\mathcal{N}_{i}}\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
where $\mathcal{P}$ denotes the set of positive node pairs (e.g., semantically or spatially close), $\mathcal{N}_{i}$ includes both positives and negatives for node $i$, and $\tau$ is a temperature hyperparameter controlling the sharpness of the distribution. Contrastive loss enforces local consistency in the learned embeddings, ensuring that nodes with similar contexts (semantic or spatial) are embedded close together. This regularizes the representation space and improves generalization, especially in heterogeneous graphs.
@@
Alignment Loss: The alignment loss encourages node embeddings to remain close to the centroid of their assigned cluster. This constraint stabilizes training and reduces noise by enforcing tighter adherence to cluster prototypes. The loss is defined as:
$\mathcal{L}_{\text{align}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{c}_{C(i)}\right\|_{2}^{2}$
where $\mathbf{z}_{i}$ is the embedding of node $i$, and $\mathbf{c}_{C(i)}$ is the centroid of cluster $C(i)$, computed as the mean embedding of all nodes in that cluster. This loss helps to maintain cluster cohesion by anchoring each node close to its corresponding cluster center. When combined with coherence and contrastive losses, alignment loss enhances the structural integrity of the latent space, especially in high-noise environments or early training stages.
Total Loss Objective: To effectively learn cluster-aware and structure-preserving embeddings, we combine all three objectives into a unified loss function. The total loss is defined as:
@@
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{contrast}}+\beta\cdot\mathcal{L}_{\text{coherence}}+\gamma\cdot\mathcal{L}_{\text{align}}$
where $\alpha$, $\beta$, and $\gamma$ are weighting coefficients that control the relative contribution of each component. This composite formulation allows the model to jointly optimize for local consistency, global cluster structure, and prototype alignment.
@@ 3.3 Embedded Features Clustering
After obtaining the embedded features produced by the final layer of the GCN encoder, we apply unsupervised clustering to the resulting embeddings $\mathbf{z}_{i}$. Clustering helps to reveal latent topic or community structures in the joint semantic-geographic representation. We use Spectral Clustering, which first computes a similarity matrix that reflects the pairwise relationships between data points, and then utilizes the eigenvectors of the graph Laplacian to partition the data. This method is particularly effective for high-dimensional and non-convex data distributions, such as those found in graph-based embeddings [von2007]. Likewise, we fix the number of clusters $K=10$ to stabilize the process and reduce variability, ensuring consistent topic identification across runs. This choice also enhances reproducibility by maintaining the same clustering configuration.
4 Validation
A comprehensive validation was conducted to rigorously assess the performance and generalization capability of the proposed methodology. We employed a multi-faceted experimental procedure for this purpose. First, we evaluated the models’ ability to capture coherent patterns and embed the social media data into clusters that are cohesive and distinct, corresponding to high intra-cluster similarity and low inter-cluster similarity. Next, we benchmarked their performance against the most comparable method identified in the literature, as well as a sequential baseline, across four diverse datasets. Finally, the utility of the proposed models was demonstrated through a case study on the 2017 Hurricane Harvey in the United States, in which we evaluated the interpretability and the spatial autocorrelation patterns found in the relevant topics, which are proven to be significant for disaster management applications.
4.1 Experimental Setup
@@
Datasets: As the models were designed specifically for the analysis of social media posts in disaster situations, we validated our proposed models using four tweet datasets related to distinct real-world disaster scenarios: (1) the 2014 Napa earthquake [brocher2015], which occurred on August 24, 2014. The geographical bounds focus on the Napa Valley region in California, covering 4,287 km². (2) Hurricane Harvey in 2017 [noaa2017], a Category 4 hurricane that impacted the southeastern Texas region from August 25 to September 7, 2017. The dataset reflects a large spatial extent of 26,736 km², particularly around Houston. (3) The 2021 Ahr Valley flood [koks2021], caused by heavy rainfall, which took place in July 2021. The dataset covers the 1,256 km² area of the Ahr Valley in Germany. (4) The 2023 Chile wildfires [cordero2024], which affected central and southern Chile and were most intense in February 2023. The dataset spans a vast area of 546,905 km². All the datasets were collected using Twitter’s former v1.1 streaming and recent search API endpoints, while filtering by the bounding box Polygon (Geographic Bounds), and the posting time (Time Frame), as depicted in Table 1.
As a common practice in natural language processing (NLP), we preprocessed the raw text data (tweets) before performing further analysis. This preprocessing is done to ensure consistency, remove noise, and improve the quality of the data. Specifically, we removed URLs using a regular expression and converted the text to lowercase for keyword extraction. HTML entities were unescaped, Mentions (i.e., @references) were removed, and extra whitespace was stripped. Non-alphanumeric characters were removed, followed by the elimination of numbers. For BERT-specific cleaning, Mentions were replaced with @user, URLs were converted to http, and newline characters were removed. These steps ensure that the tweets are appropriately formatted for further analysis and model training.
Table 1: Properties of the datasets used for validation. The time frame and geographic bounds were used for filtering. In addition, we computed the number of posts within each dataset and the size of the cover area.
@@
Ahr Valley floods 11,177 2021-07-01 – 2021-07-31 (50.01, 6.00, 50.75, 8.00) 1,256 km²
Hurricane Harvey 28,933 2017-08-25 – 2017-09-07 (27.58, –98.77, 30.66, –90.86) 26,736 km²
Chile wildfires 100,707 2023-02-01 – 2023-02-28 (–55.95, –76.89, –17.34, –65.22) 546,905 km²
Napa earthquake 90,907 2014-08-24 – 2014-08-25 (37.11, –123.05, 38.98, –121.04) 4,287 km²

### v2

@@ 3.1 Network Architectures
A schematic overview of both architectures is provided in Fig. 1. We employ Graph Convolutional Networks (GCNs) to model local neighborhood interactions induced by semantic and geographic graph structures. GCNs operate via a localized message-passing scheme, updating node features according to
@@
where $\hat{A}$ is the normalized adjacency matrix, $\mathbf{X}^{(l)}$ denotes the feature matrix at layer $l$, $\mathbf{W}^{(l)}$ is a learnable weight matrix, and $\sigma$ is a non-linear activation function such as ReLU. In both architectures, we use two GCN layers with ReLU activations. In the MultiGraph setup, dropout and layer normalization are applied between layers to improve generalization and training stability.
In the MonoGraph architecture, the unified graph is processed by a single shared GCN encoder, yielding joint node embeddings that integrate semantic similarity and geographic proximity through shared message passing. In contrast, the MultiGraph architecture employs two separate GCN branches, one operating on the semantic graph and one on the geographic graph, producing semantic embeddings $\mathbf{x}_{i}^{(s)}$ and geographic embeddings $\mathbf{x}_{i}^{(g)}$ for each node $i$. These modality-specific embeddings are subsequently fused through a learnable multi-head cross-attention mechanism. In this formulation, geographic embeddings serve as the queries $(Q)$, while semantic embeddings act as both the keys $(K)$ and values $(V)$. Since attention is applied independently at the node level rather than over sequences, this mechanism functions as a geo-conditioned feature gating operation, allowing the model to selectively reweight semantic feature dimensions based on geographic context rather than modeling interactions across nodes. This asymmetric design is intentional: geographic context provides a stable, low-dimensional conditioning signal that modulates which semantic features are emphasized, rather than being transformed itself. Treating geographic embeddings as queries allows the model to adaptively select and reweight semantic dimensions that are most relevant for a given spatial context, while preserving the expressive capacity of the semantic representations. Formally, the fused representation is obtained as:
$\mathbf{z}_{i}=\mathrm{Attn}\bigl(\mathbf{x}_{i}^{(g)},\mathbf{x}_{i}^{(s)},\mathbf{x}_{i}^{(s)}\bigr),$
where $\mathrm{Attn}(\cdot)$ denotes multi-head scaled dot-product attention. The multi-head attention layer first linearly projects $(Q,K,V)$ into $h$ subspaces (attention heads) of dimension $d_{k}=\frac{\text{embed\_dim}}{h}$. For each head, scaled dot-product attention is computed as:
$\mathrm{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{softmax}\!\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$
and the outputs of all heads are concatenated and linearly transformed to produce the final fused embedding.
@@ 3.2 Loss Design and Training Objective
To guide the learning of node representations, we optimize a composite objective that combines contrastive, coherence, and alignment losses. The contrastive loss preserves local semantic neighborhood structure induced by the graph, while the coherence and alignment losses introduce global structural regularization by encouraging compact and well-separated clusters in the embedding space. Joint optimization of these objectives enables the model to learn representations that capture both local relational consistency and global cluster organization. During training, clustering is incorporated to provide coarse-grained global guidance that complements the local nature of contrastive learning. At each training epoch, node embeddings are clustered using the K-means algorithm with a fixed number of clusters ($k_{\text{means}}=15$) and a fixed random seed ($42$) to ensure consistency across epochs. The resulting cluster assignments are used exclusively to compute the coherence and alignment losses and act as dynamic pseudo-labels, while the contrastive loss operates solely on local graph neighborhoods. Since clustering is non-differentiable and performed outside the computational graph, it does not impose hard constraints on the embedding geometry. For evaluation and qualitative analysis, clustering is performed post hoc on the final learned embeddings using Spectral Clustering, as described in Section 3.3.
@@
Coherence Loss: The coherence loss consists of two complementary components: intra-cluster similarity, which promotes internal consistency within clusters, and inter-cluster similarity, which discourages overlap between different clusters.
The intra-cluster similarity measures how similar the embeddings of nodes within the same cluster are to each other:
$\mu_{\text{intra}}=\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}$,
@@
where $K$ is the number of clusters and $C_{k}$ denotes the set of nodes assigned to cluster $k$. Cosine similarity is computed as
$\cos(\mathbf{z}_{i},\mathbf{z}_{j})=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|_{2}\|\mathbf{z}_{j}\|_{2}}.$
Maximizing intra-cluster similarity encourages compact and internally consistent clusters in the learned embedding space.
The inter-cluster similarity captures the similarity between embeddings belonging to different clusters:
$\mu_{\text{inter}}=\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}},$
where $P_{\text{inter}}$ denotes all node pairs drawn from distinct clusters. Lower values indicate better separation between clusters. To jointly encourage intra-cluster compactness and inter-cluster separation, the coherence loss is defined as:
@@
$\mathcal{L}_{\text{coherence}}=-\left(\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}-\lambda\cdot\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}\right),$
where $\lambda$ controls the trade-off between the two terms.
@@
Contrastive Loss: To preserve local semantic or structural neighborhoods, we employ a contrastive loss that pulls embeddings of similar nodes closer while pushing dissimilar ones apart. Positive pairs are constructed from the semantic graph: for each node $i$, its positive set consists of its top-$k$ semantic neighbors connected to $i$ by edges in the semantic graph, while all other nodes within the same mini-batch act as implicit negatives. The contrastive objective is defined as:
$\mathcal{L}_{\text{contrast}}=-\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{(i,k)\in\mathcal{N}_{i}}\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)},$
where $\mathcal{P}$ denotes the set of positive node pairs, $\mathcal{N}_{i}$ contains both positive and negative samples for node $i$, and $\tau$ is a temperature parameter controlling the sharpness of the similarity distribution. Unless otherwise stated, we use a fixed temperature parameter $\tau=0.5$ during training, as identified by the sensitivity analysis in Section 6.
@@
Alignment Loss: The alignment loss encourages node embeddings to remain close to the centroid of their assigned cluster:
$\mathcal{L}_{\text{align}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{c}_{C(i)}\right\|_{2}^{2},$
where $\mathbf{c}_{C(i)}$ denotes the centroid of cluster $C(i)$, computed as the mean embedding of all nodes assigned to that cluster.
Total Loss Objective: The final training objective combines all components as:
@@
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{contrast}}+\beta\cdot\mathcal{L}_{\text{coherence}}+\gamma\cdot\mathcal{L}_{\text{align}},$
where $\alpha$, $\beta$, and $\gamma$ control the relative contribution of each loss term. Unless otherwise stated, we use the baseline configuration $\alpha=0.8$, $\beta=0.2$, $\gamma=0.1$, $k_{\text{means}}=15$, $seed=42$, and $\tau=0.5$, during training, as identified by the sensitivity analysis in this section.
@@ 3.3 Embedded Features Clustering
After obtaining the embedded features produced by the final layer of the GCN encoder, we apply unsupervised clustering to the resulting embeddings $\mathbf{z}_{i}$. Clustering helps to reveal latent topic or community structures in the joint semantic-geographic representation. We use Spectral Clustering, which first computes a similarity matrix that reflects the pairwise relationships between data points, and then utilizes the eigenvectors of the graph Laplacian to partition the data. This method is particularly effective for high-dimensional and non-convex data distributions, such as those found in graph-based embeddings [von2007]. We fix the number of clusters ($k_{\text{spect}}=10$) for post hoc analysis to stabilize the clustering process and reduce variability, ensuring consistent topic identification across runs. This choice is made independently of the training-time clustering used for representation learning and reflects a trade-off between topic granularity and interpretability commonly adopted in topic modeling evaluation. Fixing $k_{\text{spect}}$ also enhances reproducibility by maintaining a consistent clustering configuration.
4 Experimental Setup
This section describes the datasets, preprocessing steps, model configuration, and training protocol used throughout the experimental evaluation. All design choices are reported once here to ensure consistency and reproducibility across subsequent analyses, including clustering evaluation and Topic Quality (TQ) benchmarking.
4.1 Datasets
@@
As the models are designed specifically for the analysis of social media posts in disaster situations, we validated our proposed methods using four tweet datasets corresponding to distinct real-world disaster events: (1) the 2014 Napa earthquake [brocher2015], which occurred on August 24, 2014. The geographic bounds focus on the Napa Valley region in California, covering 4,287 km2; (2) Hurricane Harvey in 2017 [noaa2017], a Category 4 hurricane impacting southeastern Texas between August 25 and September 7, 2017, covering a spatial extent of 26,736 km2; (3) the 2021 Ahr Valley flood [koks2021], which occurred in July 2021 and affected a 1,256 km2 region in Germany; and (4) the 2023 Chile wildfires [cordero2024], which impacted central and southern Chile in February 2023 and span a large area of 546,905 km2. All datasets were collected using Twitter’s former v1.1 streaming and recent search API endpoints, filtered by geographic bounding polygons and event-specific time frames  [^1] The data supporting the findings of this study are available from the authors upon reasonable request, subject to platform terms and conditions. . Table 1 summarizes the key properties of each dataset.
Table 1: Properties of the datasets used for validation.
@@
Ahr Valley floods 11,177 2021-07-01 – 2021-07-31 (50.01, 6.00, 50.75, 8.00) 1,256 km2
Hurricane Harvey 28,933 2017-08-25 – 2017-09-07 (27.58, –98.77, 30.66, –90.86) 26,736 km2
Chile wildfires 100,707 2023-02-01 – 2023-02-28 (–55.95, –76.89, –17.34, –65.22) 546,905 km2
Napa earthquake 90,907 2014-08-24 – 2014-08-25 (37.11, –123.05, 38.98, –121.04) 4,287 km2
4.2 Preprocessing
As is standard practice in natural language processing, the raw tweet text was preprocessed prior to analysis to ensure consistency and reduce noise. URLs were removed using regular expressions and all text was lowercased. HTML entities were unescaped, user mentions were removed, and excess whitespace was stripped. Non-alphanumeric characters and numeric tokens were eliminated. For embedding generation, we follow the input preprocessing conventions recommended for Sentence-BERT, replacing mentions with the token @user, mapping URLs to http, and removing newline characters. These steps ensure compatibility with the pre-trained language models and consistency across datasets.
4.3 Model Configuration and Training